OpenStudy (anonymous):

cos(2Tan-1(4/3)) evaluate

OpenStudy (anonymous):

is that \[\cos (2\tan^{-1}\frac{4}{3})\]

OpenStudy (anonymous):

or \[\cos ^2(\tan^{-1}\frac{4}{3})\]